Hi,I have a problem not sure how to approach it.
4s-7+=-2(1-s)

Answer by Nate(3500)

I will help you solve to find what s equals.
4s-7=-2(1-s)--original equation
4s-7=-2+2s--you'ld distribute the -2 with 1-s
2s=5--now, i got all the variables and constants on their respective sides
s=5/2--last, divide by two to get the variable, s, by itself